# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v2

@@ 1 Introduction
Concept-based Models (CBMs) are a broad class of self-explainable classifiers [1, 2, 3, 4, 5, 6, 7] designed for high performance and ante-hoc interpretability. Learning a CBM involves solving two conjoint problems: acquiring high-level concepts describing the input (e.g., an image) and an inference layer that predicts a label from them. In many applications, it is essential that these two elements are “high quality”, in the sense that: i) the concepts should be interpretable, as failure in doing so compromises understanding [8, 9] and steerability [10, 11], both key selling points of CBMs; and ii) the concepts and inference layer should behave well also out of distribution (OOD), e.g., they should not pick up spurious correlations between the input, the concepts and the output [12, 13, 14].
@@ 2 Preliminaries
A key issue is how to ensure the concepts are interpretable. Some CBMs rely on concept annotations [3, 27, 4, 5, 28, 29]. These are however expensive to obtain, prompting researchers to replace them with (potentially unreliable [30, 31]) annotations obtained from foundation models [32, 33, 34, 35] or unsupervised concept discovery [1, 2, 6, 36].
@@ 2.1 Reasoning Shortcuts
RSs by definition yield good labels in-distribution, yet they compromise out-of-distribution (OOD) performance. For instance, in autonomous driving tasks, NeSy-CBMs can confuse the concepts of “pedestrian” and “red light”, leading to poor decisions for OOD decisions where the distinction matters [22]. The meaning of concepts affected by RSs is unclear, affecting understanding [8], intervenability [54, 55, 56], debugging [57, 14] and down-stream applications that hinge on concepts being high-quality, like NeSy formal verification [58, 59, 60]. Unfortunately, existing works on RSs do not apply to CBMs and NeSy-CBMs where the inference layer is learned.
@@ 4 Practical Solutions
Supervised strategies. The most direct strategies for controlling the semantics of learned CBMs rely on supervision. Concept supervision is the go-to solution for improving concept quality in CBMs [3, 71, 4, 5] and NeSy-CBMs [22, 21]. Full concept supervision prevents learning $\bm{\mathrm{\alpha}}\neq\mathrm{id}$. However, this does not translate into guarantees on the inference layer $\bm{\mathrm{\beta}}$, at least in tasks affected by confounding factors such as spurious correlations among concepts [14]. E.g., if the $i$-th and $j$-th ground-truth concepts are strongly correlated, the inference layer $\bm{\mathrm{\beta}}$ can interchangeably use either, regardless of what $\bm{\mathrm{\beta}}^{*}$ does.
@@
Both supervised [5] and unsupervised [1, 76] CBMs may employ a reconstruction penalty [77, 78] to encourage learning informative concepts. Reconstruction can help prevent collapsing distinct ground-truth concepts into single learned concepts, e.g., in MNIST-SumParity the odd digits cannot be collapsed together without impairing reconstruction. [^3] This is provably the case context-style separation i.e., assuming concepts are independent of stylistic features like calligraphy [19, Proposition 6]. Alternatively, one can employ entropy maximization [79] to spread concept activations evenly across the bottleneck. This can be viewed as a less targeted but more efficient alternative to reconstruction, which becomes impractical for higher dimensional inputs. Another option is contrastive learning [80], in that augmentations render learned concepts more robust to style variations [81], e.g., for MNIST-based tasks it helps to cluster distinct digits [82]. Unsupervised strategies all implicitly counteract simplicity bias whereby $\bm{\mathrm{\alpha}}$ ends up collapsing.
@@ 5 Case Studies
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [41, 79] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [3] with no supervision on the concepts. Deep Symbolic Learning (DSL) [50] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [63], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [1], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [3] using only label supervision and do not rely on foundation models for gathering annotations [32, 33, 34]. For additional details, see Appendix˜A.
@@ 6 Related Work
Shortcuts in machine learning. Shortcuts occur when machine learning models solve a prediction task by using features that correlate with but do not cause the desired output, leading to poor OOD behavior [12, 10, 85, 86]. Existing work focuses on black-box models rather than CBMs. Existing studies on the semantics of CBM concepts [87, 15, 88, 89] focus on individual failures but lack formal notion of intended semantics. One exception is [17] which, however, ignores the role of the inference layer. We build on known results on reasoning shortcuts [90, 19, 20, 53] which are restricted to NeSy-CBMs in which the prior knowledge is given and fixed. Our analysis upgrades these results to general CBMs. Furthermore, our simulations indicate that strategies that are effective for CBMs and RSs no longer work for JRSs. At the same time, while NeSy approaches that learn prior knowledge and concepts jointly are increasingly popular [48, 91, 49, 50, 51, 52], existing work neglects the issue of shortcuts in this setting. Our work partially fills this gap. Shortcuts and JRSs are related but distinct. While shortcuts can induce JRSs, the converse is not true: a model can be affected by a JRS even if it does not rely on confounders, as shown in Fig.˜2. This distinction also implies that mitigation strategies designed for vanilla shortcuts do not necessarily transfer to JRSs.
Relation to identifiability. Works in Independent Component Analysis and Causal Representation Learning focus on the identifiability of representations up to an equivalence relation, typically up to permutations and rescaling [92, 93, 69] or more general transformations [94, 95]. The equivalence relation introduced along with intended semantics (Definition˜3.3) establishes a specific connection between the maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. This aligns with existing works that aim to identify representations and linear inference layers using supervised labels [68, 96, 97, 98, 99]. Moreover, while prior works primarily focus on identifying real-valued representations, we examine the identifiability of discrete concepts and of the inference layer. [100] explores the identifiability of (potentially discrete) concepts linearly encoded in model representations but focuses on multi-environment settings. Our results differ in that we show how concepts can be identified in CBMs by circumventing reasoning shortcuts.
@@ 7 Conclusion
We study the issue of learning CBMs and NeSy-CBMs with high-quality concepts and inference layers. We formalize intended semantics and joint reasoning shortcuts (JRSs), showing how, under suitable assumptions, zeroing the number of deterministic JRSs provably prevents all JRSs, yielding identifiability. Numerical simulations indicate that JRSs are very frequent and impactful, and that the only (partially) beneficial mitigation stratgy is contrastive learning. Our work paves the way to the design of effective solutions and therefore more robust and interpretable CBMs. In future work, we plan to extend our theory to general neural nets and large language models, and we will take a closer look at more complex mitigations such as smartly constraining the inference layer [49, 52], debiasing [13], and human interaction [10]. Another asset is considering concept-level supervision gathered from foundation models like CLIP, however recent results [35] call for caution in using them to supervise the bottleneck. Wrong concept-level supervision may transfer biases of foundation models and induce JRSs, thus resulting ineffective as a mitigation strategy.
@@ A.4 Clevr
Clevr [83] consists of of 3D scenes of simple objects rendered with Blender, see Fig.˜4. Images are $3\times 128\times 128$ and contain a variable number of objects. We consider only images with $2$ to $4$ objects each, primarily due to scalability issues with DPL. The ground-truth labels were computed by applying the rules (prior knowledge) proposed by [14], reported in Table˜8. Objects are entirely determined by four concepts, namely shape, color, material, and size. We list all possible values in Table˜7. Overall, Clevr has $6000$ training examples, $1200$ validation examples, and $1800$ test examples.
@@ A.6 Metrics
To compute the Concept Collapse metric $\mathsf{Cls}(C)$ [22], we follow the procedure described in Bortolotti et al. [22], which we briefly describe here. Specifically, we extract a confusion matrix in the multilabel setting by encoding each binary concept vector (e.g., $(0,1,1)\mapsto 3$) into a categorical label. Let $\mathcal{C}^{*}\subseteq\{0,1\}^{k}$ be the set of ground-truth concept vectors in the test set, and $\mathcal{C}$ be the set of predicted concept vectors. After converting both sets to categorical values, denoted $\mathcal{F}(\mathcal{C}^{*})$ and $\mathcal{F}(\mathcal{C})$ respectively, we construct the confusion matrix $C\in[0,1]^{m\times m}$, where $m=|\mathcal{F}(\mathcal{C})\cup\mathcal{F}(\mathcal{C}^{*})|$. From this matrix, we compute the number of predicted categories $p$ as:
@@ A.7 Architectures
Details of SENN. SENN [1] are unsupervised concept-based models consisting of a neural network that produces a distribution over concepts, a neural network that assigns a score to those concepts, and a decoder. The final prediction is made by computing the dot product between the predicted concepts and their predicted scores. SENN by default includes a reconstruction penalty, a robustness loss, and a sparsity term.
@@ D.4 Reconstruction penalty
The proof for this can be found in [19, Proposition 6]. Notice that, with full support over $\mathcal{G}$, the only such maps $\bm{\mathrm{\alpha}}$ must not confuse one concept vector for another, i.e., at most there are at most $|\mathcal{G}|!$ different valid $\bm{\mathrm{\alpha}}$’s that are essentially a bijection from $\mathcal{G}$ to $\mathcal{C}$.

### v3

@@ 1 Introduction
Concept-based Models (CBMs) are a broad class of self-explainable classifiers [alvarez2018towards, 10, 39, 21, 54, 78, 63] designed for high performance and ante-hoc interpretability. Learning a CBM involves solving two conjoint problems: acquiring high-level concepts describing the input (e.g., an image) and an inference layer that predicts a label from them. In many applications, it is essential that these two elements are “high quality”, in the sense that: i) the concepts should be interpretable, as failure in doing so compromises understanding [71, 62] and steerability [teso2023leveraging, 28], both key selling points of CBMs; and ii) the concepts and inference layer should behave well also out of distribution (OOD), e.g., they should not pick up spurious correlations between the input, the concepts and the output [geirhos2020shortcut, bahadori2021debiasing, stammer2021right].
@@ 2 Preliminaries
A key issue is how to ensure the concepts are interpretable. Some CBMs rely on concept annotations [39, 69, 21, 54, 36, 16]. These are however expensive to obtain, prompting researchers to replace them with (potentially unreliable [31, 77]) annotations obtained from foundation models [61, 93, 74, 15] or unsupervised concept discovery [alvarez2018towards, 10, 78, 70].
@@ 2.1 Reasoning Shortcuts
RSs by definition yield good labels in-distribution, yet they compromise out-of-distribution (OOD) performance. For instance, in autonomous driving tasks, NeSy-CBMs can confuse the concepts of “pedestrian” and “red light”, leading to poor decisions for OOD decisions where the distinction matters [4]. The meaning of concepts affected by RSs is unclear, affecting understanding [71], intervenability [72, 22, 76], debugging [lertvittayakumjorn2020find, stammer2021right] and down-stream applications that hinge on concepts being high-quality, like NeSy formal verification [88, 98, 59]. Unfortunately, existing works on RSs do not apply to CBMs and NeSy-CBMs where the inference layer is learned.
@@ 4 Practical Solutions
Supervised strategies. The most direct strategies for controlling the semantics of learned CBMs rely on supervision. Concept supervision is the go-to solution for improving concept quality in CBMs [39, 12, 21, 54] and NeSy-CBMs [4, 91]. Full concept supervision prevents learning $\bm{\mathrm{\alpha}}\neq\mathrm{id}$. However, this does not translate into guarantees on the inference layer $\bm{\mathrm{\beta}}$, at least in tasks affected by confounding factors such as spurious correlations among concepts [stammer2021right]. E.g., if the $i$-th and $j$-th ground-truth concepts are strongly correlated, the inference layer $\bm{\mathrm{\beta}}$ can interchangeably use either, regardless of what $\bm{\mathrm{\beta}}^{*}$ does.
@@
Both supervised [54] and unsupervised [alvarez2018towards, 44] CBMs may employ a reconstruction penalty [2, 37] to encourage learning informative concepts. Reconstruction can help prevent collapsing distinct ground-truth concepts into single learned concepts, e.g., in MNIST-SumParity the odd digits cannot be collapsed together without impairing reconstruction. [^3] This is provably the case context-style separation i.e., assuming concepts are independent of stylistic features like calligraphy [56, Proposition 6]. Alternatively, one can employ entropy maximization [51] to spread concept activations evenly across the bottleneck. This can be viewed as a less targeted but more efficient alternative to reconstruction, which becomes impractical for higher dimensional inputs. Another option is contrastive learning [11], in that augmentations render learned concepts more robust to style variations [83], e.g., for MNIST-based tasks it helps to cluster distinct digits [68]. Unsupervised strategies all implicitly counteract simplicity bias whereby $\bm{\mathrm{\alpha}}$ ends up collapsing.
@@ 5 Case Studies
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [50, 51] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [39] with no supervision on the concepts. Deep Symbolic Learning (DSL) [13] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [53], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [alvarez2018towards], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [39] using only label supervision and do not rely on foundation models for gathering annotations [61, 93, 74]. For additional details, see Appendix˜A.
@@ 6 Related Work
Shortcuts in machine learning. Shortcuts occur when machine learning models solve a prediction task by using features that correlate with but do not cause the desired output, leading to poor OOD behavior [geirhos2020shortcut, teso2023leveraging, 95, 75]. Existing work focuses on black-box models rather than CBMs. Existing studies on the semantics of CBM concepts [57, 49, 29, 65] focus on individual failures but lack formal notion of intended semantics. One exception is [55] which, however, ignores the role of the inference layer. We build on known results on reasoning shortcuts [45, 56, 84, 80] which are restricted to NeSy-CBMs in which the prior knowledge is given and fixed. Our analysis upgrades these results to general CBMs. Furthermore, our simulations indicate that strategies that are effective for CBMs and RSs no longer work for JRSs. At the same time, while NeSy approaches that learn prior knowledge and concepts jointly are increasingly popular [85, 94, 48, 13, 79, 87], existing work neglects the issue of shortcuts in this setting. Our work partially fills this gap. Shortcuts and JRSs are related but distinct. While shortcuts can induce JRSs, the converse is not true: a model can be affected by a JRS even if it does not rely on confounders, as shown in Fig.˜2. This distinction also implies that mitigation strategies designed for vanilla shortcuts do not necessarily transfer to JRSs.
Relation to identifiability. Works in Independent Component Analysis and Causal Representation Learning focus on the identifiability of representations up to an equivalence relation, typically up to permutations and rescaling [34, 26, 82] or more general transformations [67, 5]. The equivalence relation introduced along with intended semantics (Definition˜3.3) establishes a specific connection between the maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. This aligns with existing works that aim to identify representations and linear inference layers using supervised labels [41, 23, 3, nielsen2024challenges, nielsen2025does]. Moreover, while prior works primarily focus on identifying real-valued representations, we examine the identifiability of discrete concepts and of the inference layer. [64] explores the identifiability of (potentially discrete) concepts linearly encoded in model representations but focuses on multi-environment settings. Our results differ in that we show how concepts can be identified in CBMs by circumventing reasoning shortcuts. A concurrent work by Goyal et al. [25] identifies latent concepts through label supervision and a sparsity regularization, showing that a variant of CBNMs can provably recover concepts with the intended semantics and thus avoiding JRSs altogether, see [25, Theorem 2].
@@ 7 Conclusion
We study the issue of learning CBMs and NeSy-CBMs with high-quality concepts and inference layers. We formalize intended semantics and joint reasoning shortcuts (JRSs), showing how, under suitable assumptions, zeroing the number of deterministic JRSs provably prevents all JRSs, yielding identifiability. Numerical simulations indicate that JRSs are very frequent and impactful, and that the only (partially) beneficial mitigation stratgy is contrastive learning. Our work paves the way to the design of effective solutions and therefore more robust and interpretable CBMs. In future work, we plan to extend our theory to general neural nets and large language models, and we will take a closer look at more complex mitigations such as smartly constraining the inference layer [48, 87], debiasing [bahadori2021debiasing], and human interaction [teso2023leveraging]. Another asset is considering concept-level supervision gathered from foundation models like CLIP, however recent results [15] call for caution in using them to supervise the bottleneck. Wrong concept-level supervision may transfer biases of foundation models and induce JRSs, thus resulting ineffective as a mitigation strategy.
@@ A.4 Clevr
Clevr [33] consists of of 3D scenes of simple objects rendered with Blender, see Fig.˜4. Images are $3\times 128\times 128$ and contain a variable number of objects. We consider only images with $2$ to $4$ objects each, primarily due to scalability issues with DPL. The ground-truth labels were computed by applying the rules (prior knowledge) proposed by [stammer2021right], reported in Table˜8. Objects are entirely determined by four concepts, namely shape, color, material, and size. We list all possible values in Table˜7. Overall, Clevr has $6000$ training examples, $1200$ validation examples, and $1800$ test examples.
@@ A.6 Metrics
To compute the Concept Collapse metric $\mathsf{Cls}(C)$ [4], we follow the procedure described in Bortolotti and others [4], which we briefly describe here. Specifically, we extract a confusion matrix in the multilabel setting by encoding each binary concept vector (e.g., $(0,1,1)\mapsto 3$) into a categorical label. Let $\mathcal{C}^{*}\subseteq\{0,1\}^{k}$ be the set of ground-truth concept vectors in the test set, and $\mathcal{C}$ be the set of predicted concept vectors. After converting both sets to categorical values, denoted $\mathcal{F}(\mathcal{C}^{*})$ and $\mathcal{F}(\mathcal{C})$ respectively, we construct the confusion matrix $C\in[0,1]^{m\times m}$, where $m=|\mathcal{F}(\mathcal{C})\cup\mathcal{F}(\mathcal{C}^{*})|$. From this matrix, we compute the number of predicted categories $p$ as:
@@ A.7 Architectures
Details of SENN. SENN [alvarez2018towards] are unsupervised concept-based models consisting of a neural network that produces a distribution over concepts, a neural network that assigns a score to those concepts, and a decoder. The final prediction is made by computing the dot product between the predicted concepts and their predicted scores. SENN by default includes a reconstruction penalty, a robustness loss, and a sparsity term.
@@ D.4 Reconstruction penalty
The proof for this can be found in [56, Proposition 6]. Notice that, with full support over $\mathcal{G}$, the only such maps $\bm{\mathrm{\alpha}}$ must not confuse one concept vector for another, i.e., at most there are at most $|\mathcal{G}|!$ different valid $\bm{\mathrm{\alpha}}$’s that are essentially a bijection from $\mathcal{G}$ to $\mathcal{C}$.